OpenStudy (ledah):

simplify sqrt(25)z^4

OpenStudy (ledah):

\[\sqrt{25z^4}\]

OpenStudy (anonymous):

sqrt(ab) = sqrt(a)sqrt(b)

OpenStudy (ledah):

Not sure I follow

OpenStudy (anonymous):

Well, take a as 25 and b as z^4

OpenStudy (ledah):

\[5\sqrt{z^4}\]

OpenStudy (anonymous):

Ok, you square rooted the 25 , what about the other bit?

OpenStudy (ledah):

would it look lke this \[5z ^{2}\]

OpenStudy (anonymous):

It would indeed.....:-)

OpenStudy (ledah):

okay thank you for the help

OpenStudy (anonymous):

ur welcome